### v3

# Heat dissipation in marginally stable linear time-delayed Langevin systems

Abstract
The thermodynamic properties of time-delayed dynamics remain largely unexplored, especially for systems that exhibit asymptotically non-stationary behavior. Here, we investigate heat dissipation in two classes of marginally stable linear time-delayed Langevin dynamics: (i) diffusive criticality, which asymptotically manifests as scaled Brownian diffusion, and (ii) oscillatory criticality, which shows oscillation with diffusive amplitude. By analytical derivations, we find fundamentally different thermodynamic signatures: the average heat dissipation rate asymptotically approaches a constant for diffusive criticality but diverges linearly with oscillations for oscillatory criticality, despite both showing linearly growing variance over time. We further numerically study the probability distributions of heat dissipation rates for both types of critical dynamics. Our results demonstrate that non-stationary time-delayed dynamics with similar scaling can yield qualitatively distinct heat dissipation behaviors, depending on the underlying dynamical details. This work provides a concrete foundation for future investigations into thermodynamic properties of general nonlinear time-delayed systems.
Introduction
Time-delayed systems are ubiquitous in natural and engineered systems [1, 2, 3, 4, 5, 6, 7], where finite signal propagation or processing time introduces nontrivial memory effects. Time-delayed effects may induce complex phenomena, including spontaneous oscillations [8, 7, 9], bifurcations [10, 11], resonance [12, 13, 14], chaos [15], synchronization [16, 17, 18], and pattern formation [19, 20]. Investigating these delay-induced effects is crucial for the accurate analysis and effective design of real-world systems.
Time-delayed systems can also manifest nontrivial thermodynamic properties due to their inherent non-Markovian nature. Several studies reveal non-zero (sometimes even negative) average heat dissipation in time-delayed Langevin systems in the steady state, highlighting their non-equilibrium characters [21, 22, 23, 24, 25, 26]. Thermodynamic quantities are also discussed using generalized fluctuation theorems and thermodynamic uncertainty relations adapted for time-delayed dynamics [27, 28, 29, 30, 31, 32, 33, 34]. However, current research on the thermodynamic properties of time-delayed systems has primarily focused on stable dynamics (i.e., dynamics that permits a steady state). For systems in non-stationary states, including those exhibiting diffusive or oscillatory behavior under Pyragas time-delayed feedback control [35, 36, 37], the thermodynamic properties have not been comprehensively explored. Even for linear time-delayed Langevin dynamics, a systematic study is lacking on the heat dissipation at marginal stability (when the parameters of the system lie on the boundary of the stability conditions [38]).
To address this gap, we study the heat dissipation in two classes of marginally stable linear time-delayed Langevin dynamics: (i) diffusive criticality and (ii) oscillatory criticality. We evaluate the average heat dissipation rate through the variance of the system’s position, which is asymptotically determined by the roots of the linear stability equation with zero real parts. Surprisingly, we find that although both types of criticality exhibit linearly divergent variance in time, they show fundamentally different thermodynamic behaviors: the average heat dissipation rate approaches a constant in the long-time limit for diffusive criticality but diverges linearly in time for oscillatory criticality. Moreover, we also examine the distribution of the heat dissipation via numerical simulations, which further reveal distinct statistical properties for each criticality class. Our results for these linear systems demonstrate that non-stationary time-delayed dynamics with similar scaling can produce qualitatively different heat dissipation, depending on the details of the dynamics. This study provides a concrete example describing the thermodynamic properties in time-delayed Langevin systems beyond the stable regime, laying a foundation for broader investigations into nonlinear systems.
This work is structured as follows: Sec. II introduces the linear time-delayed Langevin dynamics and specifies the two classes of marginally stable dynamics. Sec. III presents the core of our work: the analytical derivation of the asymptotic average heat dissipation rates for both critical dynamics, alongside numerical results that validate our derivation and explore the full distribution of dissipation. Finally, Sec. IV summarizes this work and discusses the future directions.
Marginally stable dynamics of linear time-delayed Langevin systems
Consider a class of time-delayed systems governed by the overdamped Langevin equation:
[…]
where $\tau>0$ is a fixed delay time, $a$ denotes the strength of a linear non-delayed force, and $b$ represents the strength of the linear time-delayed force. The friction coefficient is given by $\gamma$, and $\xi(t)$ is white Gaussian noise satisfying $\left\langle\xi(t)\right\rangle=0$ and $\left\langle\xi(t)\xi(t^{\prime})\right\rangle=2\gamma T\delta(t-t^{\prime})$, where $T$ is the temperature of the thermal environment. The dynamics is initialized at $t=t_{\text{init}}$ subjected to a predefined historical function $\phi(t)$ defined for $t\in[t_{\text{init}}-\tau,t_{\text{init}}]$. Throughout this work, we set $t_{\text{init}}=0$ and $\gamma=1$ without loss of generality.
The solution to the dynamics described by Eq. (1) is as follows [38]:
$x(t)$ $=x_{0}(t)\phi(0)+b\int_{-\tau}^{0}x_{0}(t-s-\tau)\phi(s)ds$
$\quad+\int_{0}^{t}x_{0}(t-s)\xi(s)ds,$ (2)
where the fundamental solution $x_{0}(t)$ is defined as the solution of Eq. (1) under the following conditions:
$T=0,\ \ \phi(t)=0\ (t<0),\ \ \phi(0)=1.$ (3)
In Appendix B, we show that
$x_{0}(t)=\sum_{k=-\infty}^{\infty}\dfrac{1}{1+b\tau e^{-S_{k}\tau}}e^{S_{k}t}\ \ (t>0),$ (4)
where $\{S_{k}\}\ (k\in\mathcal{Z})$ correspond to all the roots that satisfy the characteristic equation [39]
$S_{k}-a-be^{-S_{k}\tau}=0.$ (5)
This characteristic equation corresponds to the linear stability of the dynamics (Eq. (1)). The distribution of $x(t)$ is Gaussian, with the mean value and variance:
$\langle x(t)\rangle$ $=x_{0}(t)\phi(0)+b\int_{-\tau}^{0}x_{0}(t-s-\tau)\phi(s)ds,$ (6)
$\langle x^{2}(t)\rangle$ $-\langle x(t)\rangle^{2}=2T\int_{0}^{t}x_{0}^{2}(s)ds.$ (7)
Parameters $a,b,\tau$ in Eq. (5) determine the distribution of characteristic roots $\{S_{k}\}$, which further control the asymptotic behavior of the solution $x(t)$ . If $\max\{\text{Re}[S_{k}]\}<0$, the distribution of $x(t)$ becomes stationary as $t\rightarrow\infty$ and the dynamics is stable; if $\max\{\text{Re}[S_{k}]\}=0$, the variance of $x(t)$ diverges linearly in time and the dynamics is marginally stable; if $\max\{\text{Re}[S_{k}]\}>0$, the dynamics exponentially diverges in time.
There are exactly two classes of marginally stable dynamics [40, 41, 42]:
Figure 1: (a) A single trajectory simulation of the dynamics with diffusive criticality. Here a=5a=-5, $b=5$, $\tau=1$, $T=1$, and $\phi(t)=0\ \ (t\leq 0)$. (b) 100 trajectories of the dynamics with diffusive criticality for each set of parameters. Blue: $a=0.5$ and b=0.5b=-0.5; red: $a=b=0$ (Brownian motion) ; green: a=1a=-1 and $b=1$. Here $T=1$, $\tau=1$, and $\phi(t)=0\ \ (t\leq 0)$.
(i) Diffusive criticality: When
$a+b=0,\ a\tau<1,$ (8)
there is exactly one zero root in $\{S_{k}\}$ and all the others have negative real parts (we exclude $a\tau=1$ as the coefficients in Eq. (4) diverge). Such critical dynamics corresponds to scenarios of systems under Pyragas control [43, 35, 36, 37]. Under this condition (Eq. (8)), the fundamental solution satisfies
$\lim_{t\rightarrow{\infty}}x_{0}(t)=\dfrac{1}{1-a\tau}.$ (9)
Combining Eq. (9) with Eqs. (6) and (7), we obtain the dynamics in the long-time limit ($t\rightarrow\infty$)
$\langle x(t)\rangle=\frac{1}{1-a\tau}\phi(0)-\frac{a}{1-a\tau}\int_{-\tau}^{0}\phi(s)ds,\ \ (t\rightarrow\infty)$ (10)
$\langle x^{2}(t)\rangle-\langle x(t)\rangle^{2}=\frac{2T}{(1-a\tau)^{2}}\cdot t+C_{0},\ \ (t\rightarrow\infty)$ (11)
where
$C_{0}=2T\int_{0}^{\infty}\left[x_{0}^{2}(s)-\left(\dfrac{1}{1-a\tau}\right)^{2}\right]ds,$ (12)
is a bounded constant  [^1] The convergence can be verified by noticing $x_{0}(0)=\sum_{k=-\infty}^{\infty}\dfrac{1}{1+b\tau e^{-S_{k}\tau}}=1$ and accordingly $x_{0}(t)=\frac{1}{1-a\tau}+O(e^{\gamma t})$, where $\gamma$ is the largest nonzero real part of the characteristic roots. Similar discussion applies for Eq. (18). See also [53].. Note that Eq. (11) has been analytically derived in Ref. [45]. The mean displacement $\langle x(t)\rangle$ asymptotically approaches a constant that depends on the historical trajectory $\phi(t)$. The diffusion of the critical dynamics resembles free Brownian motion with a scaling factor $1/(1-a\tau)^{2}$ (Fig. 1(a)). A positive time-delayed force (b=a>0b=-a>0) suppresses the diffusion and a negative one ($b=-a<0,\ a\tau<1$) enlarges the diffusion (Fig. 1(b)).
(ii) Oscillatory criticality: When
[…]
there are exactly two conjugate imaginary roots $\pm i\sqrt{b^{2}-a^{2}}$ and all the others have negative real parts. Such oscillatory critical dynamics shows oscillation with a fluctuating amplitude (Fig. 2(a)). The fundamental solution in the long-time limit is
$\lim_{t\rightarrow\infty}x_{0}(t)$ $=\frac{1}{1+b\tau e^{-i\omega_{c}\tau}}e^{i\omega_{c}t}+\frac{1}{1+b\tau e^{i\omega_{c}\tau}}e^{-i\omega_{c}t}$
$=\dfrac{2\cos(\omega_{c}t+\theta)}{\sqrt{1+b^{2}\tau^{2}-2a\tau}},$ (14)
where
$\omega_{c}=\sqrt{b^{2}-a^{2}},\ \ \tau=\dfrac{\arccos(-\frac{a}{b})}{\sqrt{b^{2}-a^{2}}},$ (15)
$\theta=\arctan\left(\frac{\sqrt{b^{2}-a^{2}}\cdot\tau}{a\tau-1}\right).$
The mean displacement in the long-time limit ($t\rightarrow\infty$)
$\langle x(t)\rangle=\dfrac{2}{\sqrt{1+b^{2}\tau^{2}-2a\tau}}$
$\cdot\left(\phi(0)\cos(\omega_{c}t+\theta)+b\int_{-\tau}^{0}\cos[\omega_{c}(t-s-\tau)+\theta]\phi(s)ds\right)$ (16)
and the variance
$\langle x^{2}(t)\rangle-\langle x(t)\rangle^{2}=$
$\dfrac{4T\bigg{[}t+\frac{1}{2\omega_{c}}\bigg{(}\sin[2(\omega_{c}t+\theta)]-\sin(2\theta)\bigg{)}\bigg{]}}{1+b^{2}\tau^{2}-2a\tau}+C_{1},$ (17)
where
$C_{1}=2T\int_{0}^{\infty}\bigg{(}x_{0}^{2}(s)-\dfrac{4\cos^{2}(\omega_{c}s+\theta)}{1+b^{2}\tau^{2}-2a\tau}\bigg{)}ds,$ (18)
is a bounded constant. The mean displacement shows a superposition of oscillations that depends on the historical trajectory $\phi(t)$ and the variance contains a linearly increasing term and an oscillating term. For a fixed $a$, a larger strength of the delayed force $|b|$ leads to oscillations with a shorter period and an amplitude that diffuses faster (Fig. 2(b)).
Figure 2: (a) A single trajectory simulation of the dynamics with oscillatory criticality. Here a=2a=-2, b=3b=-3, $\tau\approx 1.029$, $T=1$, and $\phi(t)=0\ \ (t\leq 0)$. (b) 100 trajectories of the dynamics with oscillatory criticality for each set of parameters. Blue: a=2a=-2, b=10b=-10, and $\tau\approx 0.181$; red: a=2a=-2, b=5b=-5, and $\tau\approx 0.433$; green: a=2a=-2, b=2.5b=-2.5, and $\tau\approx 1.665$. Here $T=1$ and $\phi(t)=0\ \ (t\leq 0)$.
Heat dissipation of marginal stable dynamics
In the following, we analytically derive the asymptotic expressions for the average heat dissipation rate of two types of marginally stable dynamics, along with numerical verification. This constitutes the main result of our work.
We define the heat dissipation rate as
$dq(t)\equiv[\gamma\dot{x}(t)-\xi(t)]\circ dx(t),$ (19)
where $\circ$ denotes the Stratonovich product [46]. It represents the heat dissipated from the system into the surrounding thermal medium per unit time. The heat dissipation rate of steady state for stable solutions has been derived in [21, 22, 26]. However, those methods are not applicable when calculating the heat dissipation of marginally stable dynamics. In this section, we explicitly calculate the heat dissipation for the two classes of marginally stable dynamics.
From the definition in Eq. (19) and the dynamics described by Eq. (1), the heat dissipation of the system within an infinitesimal time interval $dt$ is given by ($\gamma\equiv 1$):
$\langle dq(t)\rangle=\left\langle\bigg{[}\frac{dx(t)}{dt}-\xi(t)\bigg{]}\circ dx(t)\right\rangle$
$=\left\langle\big{[}ax(t)+bx(t-\tau)\big{]}\circ\big{[}ax(t)+bx(t-\tau)+\xi(t)\big{]}dt\right\rangle$
$=a^{2}\langle x^{2}(t)\rangle dt+b^{2}\langle x^{2}(t-\tau)\rangle dt$
$\quad\quad+2ab\langle x(t)x(t-\tau)\rangle dt+aTdt,$ (20)
where in the last line we use
$\langle x(t)\circ\xi(t)dt\rangle=\left\langle\frac{x(t+dt)+x(t-dt)}{2}\cdot\xi(t)dt\right\rangle$
$=\frac{1}{2}\int_{0}^{t+dt}x_{0}(t-s)\langle\xi(s)ds\cdot\xi(t)dt\rangle$
$=\frac{1}{2}\left(\int_{0}^{t+dt}x_{0}(t-s)\cdot 2T\delta(t-s)ds\right)\cdot dt$
$=\frac{1}{2}\cdot x_{0}(0)\cdot 2Tdt=Tdt\quad(\text{since}\ x_{0}(0)=1)$ (21)
and
$\langle x(t-\tau)\circ\xi(t))=0$ (22)
because the system’s state at past is independent of the white noise at present. The correlation $\langle x(t)x(t-\tau)\rangle$ can be calculated through
$\frac{1}{2}d\langle x^{2}(t)\rangle=\langle x(t)\circ dx(t)\rangle$
$=a\langle x^{2}(t)\rangle dt+b\langle x(t)x(t-\tau)\rangle dt+\langle x(t)\circ\xi(t)dt\rangle,$ (23)
which gives
$b\langle x(t)x(t-\tau)\rangle=\frac{1}{2}\frac{d\langle x^{2}(t)\rangle}{dt}-a\langle x^{2}(t)\rangle-T.$ (24)
Therefore, the average heat dissipation rate is
$\left\langle\frac{dq}{dt}\right\rangle=b^{2}\langle x^{2}(t-\tau)\rangle-a^{2}\langle x^{2}(t)\rangle+a\frac{d\langle x^{2}(t)\rangle}{dt}-aT.$ (25)
Below, we discuss the heat dissipation rate of two types of marginally stable dynamics (we set $\phi(t)=0\ \ (t<0)$ for simplicity).
(i) Diffusive criticality:
Substituting Eqs. (10) and (11) into Eq. (25), we obtain
$\left\langle\frac{dq}{dt}\right\rangle=aT\cdot\dfrac{1+a\tau}{1-a\tau}\ \ (t\rightarrow\infty)_{.}$ (26)
When calculating Eq. (26), the linearly increasing second moments in Eq. (25) cancel out, resulting in a constant heat dissipation rate in the long-time limit (Fig. 3(a)). As shown in Fig. 3(b), a larger $\tau$ leads to larger heat dissipation rate. The heat dissipation can even become negative for $a<0$ and $\tau<-1/a$, despite the diffusive nature of the dynamics. For $a>0$, the heat dissipation rate diverges as $\tau$ approaches the stability boundary $a\tau=1$.
Figure 3: (a) Time evolution of the average heat dissipation rate in diffusive critical dynamics. Solid line: analytical results (calculated by Eqs. (7), (8) and  (25), where $x_{0}(t)$ is calculated through the numerical integral of Eq. (31) ); dashed line: the asymptotic behavior (Eq. (26)); crosses: numerical simulations with $10^{6}$ trajectories of precision $dt=10^{-4}$, where each point of $\Delta q/\Delta t$ is averaged over interval $\Delta t=0.1$. Here a=5a=-5, $b=5$, $\tau=1$, $T=0.1$, and $\phi(t)=0\ (t\leq 0)$. (b) Average heat dissipation rate of the diffusive critical dynamics in the long-time limit. Here $T=1$.
Figure 4: (a) Time evolution of the average heat dissipation rate in oscillatory critical dynamics. Solid line: the analytical result (calculated by Eqs. (7), (13), and (25), where $x_{0}(t)$ is calculated through the numerical integral of Eq. (31) ); dashed line: the asymptotic behavior (Eq. (27)); crosses: numerical simulations with $10^{6}$ trajectories of precision $dt=10^{-4}$, where each point of $\Delta q/\Delta t$ is averaged over interval $\Delta t=0.1$. Here a=2a=-2, b=2.1b=-2.1, $\tau\approx 4.422$, $T=0.1$, and $\phi(t)=0\ (t\leq 0)$. (b) Divergence speed of the average heat dissipation rate in oscillatory critical dynamics (Eq. (30)). Here b<|a|b<-|a| and $T=1$.
(ii) Oscillatory criticality: Substituting Eqs. (16) and (II) into Eq. (25), we obtain
$\left\langle\frac{dq}{dt}\right\rangle$ $=\frac{4(b^{2}-a^{2})T}{1+b^{2}\tau^{2}-2a\tau}\cdot t\ +A_{0}(t)+B_{0}\ \ (t\rightarrow\infty),$ (27)
where the oscillating term
[…]
and the constant term
[…]
$\ \ +(b^{2}-a^{2})C_{1}-aT.$ (29)
Note that $A_{0}(t)+B_{0}\sim O(1)$; therefore
$\lim_{t\rightarrow\infty}\frac{1}{t}\left\langle\frac{dq}{dt}\right\rangle=\frac{4(b^{2}-a^{2})T}{1+b^{2}\tau^{2}-2a\tau}.$ (30)
The heat dissipation rate of the dynamics of oscillatory criticality linearly diverges in time accompanied by an oscillation, as shown in Fig. 4(a). Its divergence rate (Eq. (30)) increases with $|b|$ for a fixed $a$, and vice versa (Fig. 4(b)). This shows a clear contrast to the heat dissipation of diffusive critical dynamics, although both dynamics show a variance linearly diverging in time.
The sharp difference in heat dissipation between the two types of criticality stems from their underlying dynamical details. The dynamics of oscillatory criticality shows oscillation with a linearly growing amplitude on average. For the majority of trajectories, the motion becomes increasingly violent as the distance covered by oscillation within a period increases with time . The resistance to the noise by such motion ends with linearly divergent heat dissipation in time. On the other hand, the dynamics of diffusive criticality manifests as a scaled Brownian motion. After a sufficiently long time, the range of diffusion remains local within a fixed interval, which leads to bounded dissipation.
(a) a=b<0a=-b<0
(b) a=b>0a=-b>0, $a\tau<1$
(c) a=2a=-2, $b=2$
(d) $a=0.5$, b=0.5b=-0.5
Figure 5: Probability distribution of the heat dissipation rate $\Delta q/\Delta t$ for diffusive critical dynamics in the long-time limit, where each sample of $\Delta q/\Delta t$ is averaged over $\Delta t=0.1$ . We take $10^{6}$ samples with simulation precision $dt=10^{-3}$ and set $T=0.1$. (a) Varying $a$ with a=b<0a=-b<0 and $\tau=1$. (b) Varying $a$ with a=b>0a=-b>0 and $\tau=1$. (c) Varying $\tau$ with a=2a=-2 and $b=2$. (d) Varying $\tau$ with $a=0.5$ and b=0.5b=-0.5.
Furthermore, we numerically examine the probability distribution of heat dissipation rate for two types of criticality. The distribution of both criticalities exhibits a single peak and an exponential asymptotic tail. For diffusive criticality, in contrast to the linearly divergent variance of the position, the distribution of heat dissipation converges to an asymptotic one in the long-time limit (Fig. 5). Larger $|a|$ and $|b|$ lead to a broader distribution of heat dissipation (Figs. 5(a) and 5(b)). The distribution also broadens as $\tau$ increases, relaxing to an asymptotic form in the large-$\tau$ limit for a=b<0a=-b<0 (Fig. 5(c)), but spreading without bound as $\tau$ approaches the stability boundary $a\tau=1$ for a=b>0a=-b>0 (Fig. 5(d)). For oscillatory criticality, the distribution continues to spread with time without reaching an asymptotic form (Fig. 6(a)). For a fixed $a$, a larger $|b|$ (and thus a smaller $\tau$ according to the marginal-stability conditions) results in a broader distribution (Figs. 6(b) and 6(c)). These features are consistent with earlier analytical results for $\left\langle dq/dt\right\rangle$, but also show how the variance of heat dissipation rate changes with the relevant parameters.
(a) a=2a=-2, b=3b=-3, $\tau\approx 2.301$
(b) a=2a=-2
(c) $a=2$
Figure 6: Probability distribution of the heat dissipation rate $\Delta q/\Delta t$ for the oscillatory critical dynamics at finite time $t$, where each sample of $\Delta q/\Delta t$ is averaged over $\Delta t=0.1$. We take $10^{6}$ samples with simulation precision $dt=10^{-3}$ and set $T=0.1$. (a) Varying the evolution time $t$ with a=2a=-2, b=3b=-3, and $\tau\approx 2.301$. (b) Varying $b$ with a=2a=-2 and $t=70$. (c) Varying $b$ with $a=2$ and $t=70$.
Discussion
In this work, we have studied the heat dissipation and its properties of two classes of criticality in marginally stable linear time-delayed Langevin systems. For diffusive criticality, despite its scaled diffusion, the probability distribution of heat dissipation becomes stationary and the average approaches a constant $aT\cdot\frac{1+a\tau}{1-a\tau}$ in the long-time limit. For oscillatory criticality, the distribution of heat dissipation continues to spread with time and the average asymptotically diverges as $\frac{4(b^{2}-a^{2})T}{1+b^{2}\tau^{2}-2a\tau}\cdot t+O(1)$ due to the diffusive amplitude of oscillation. These results show that the asymptotic behavior of the heat dissipation is determined by the characteristic roots of Eq. (5) with zero real parts for marginally stable linear time-delayed Langevin dynamics. Our results offer a reference point for further studies on stochastic thermodynamics of more general time-delayed systems, such as those under Pyragas control and near bifurcations.
One may face both significant challenges and fruitful insights when diving into the thermodynamics of the nonlinear time-delayed regime. A promising direction is to investigate stochastic resonance phenomena on nonlinear time-delayed systems. It is natural to expect that a nontrivial resonance may manifest in their heat dissipation profiles. Furthermore, nonlinearity may suppress the amplitude diffusion of bifurcations, leading to sustained finite-amplitude oscillations of time-delayed stochastic systems. Analyzing the thermodynamic properties of such oscillations would be another path for future research. Finally, the thermodynamic properties of collective behavior in time-delayed interacting stochastic systems remains an open area of research.
Acknowledgements.
X. W. thanks Naruo Ohga, Sarah A. M. Loos, and Robin A. Kopp for enlightening discussions. X. W. acknowledges financial support from Department of Physics, Graduate School of Science, the University of Tokyo.
Appendix A Numerical methods
We use the programming language Julia and the package DifferentialEquations.jl [47] and StochasticDelayDiffEq.jl [48] for numerical calculations and simulations (the Tsitouras 5/4 Runge-Kutta method with a free 4th order interpolant [49] is used for the numerical integration of $x_{0}(t)$ via Eq. (31); the Euler-Maruyama method [50] is used for trajectory simulations of stochastic dynamics). We draw the figures using the package Plots.jl [51] and GR Framework [52].
Appendix B Derivation of the exponential expansion of the fundamental solution Eq. (4)
We show how to derive Eq. (4) in detail based on the method in Ref. [39]. Consider the linear delayed differential equation (set $T=0$ in Eq. (1)):
[…]
The solution of Eq. (31) is
$x(t)=x_{0}(t)\phi(0)+b\int_{-\tau}^{0}x_{0}(t-s-\tau)\phi(s)ds,$ (32)
where the fundamental solution $x_{0}(t)$ corresponds to the solution of Eq. (31) satisfying the initial conditions $x_{0}(t)=0$ for $t<0$ and $x_{0}(0)=1$ [38].
We expand the solution of Eq. (31) using an exponential series based on the roots of characteristic equation (Eq. (5)):
$x(t)=\sum_{k=-\infty}^{\infty}C_{k}e^{S_{k}t},$ (33)
where the coefficients are left to be determined according to the dynamics ($a,b$ and $\tau$) and the initial conditions ($\phi(t)$). To determine the coefficients, we first compute the Laplace transform of Eq. (31):
$sX(s)-x(0)=aX(s)+be^{-s\tau}X(s)+b\Phi(s),$ (34)
where
$X(s)\equiv\mathcal{L}[x(t)]=\int_{0}^{\infty}e^{-st}x(t)dt,$ (35)
$\Phi(s)=\int_{0}^{\tau}e^{-st}\phi(t-\tau)dt,$ (36)
and the Laplace transform of $X(t-\tau)$ is calculated as follows
$\mathcal{L}[x(t-\tau)]=\int_{0}^{\infty}e^{-st}x(t-\tau)dt$
$=\int_{0}^{\tau}e^{-st}x(t-\tau)dt+\int_{\tau}^{\infty}e^{-st}x(t-\tau)dt$
$=\int_{0}^{\tau}e^{-st}\phi(t-\tau)dt+\int_{0}^{\infty}e^{-s(t^{\prime}+\tau)}x(t^{\prime})dt^{\prime}\quad(t^{\prime}=t-\tau)$
$=\int_{0}^{\tau}e^{-st}\phi(t-\tau)dt+e^{-s\tau}X(s)$
$=\Phi(s)+e^{-s\tau}X(s).$
From Eq. (34), we have
$X(s)=\dfrac{x(0)+b\Phi(s)}{s-a-be^{-s\tau}}.$ (37)
On the other hand, the Laplace transform of the exponential expansion representation is
$\mathcal{L}[x(t)]$ $=\mathcal{L}\left(\sum_{k=-\infty}^{\infty}C_{k}e^{S_{k}t}\right)$
$=\sum_{k=-\infty}^{\infty}\frac{C_{k}}{s-S_{k}}.$ (38)
Comparing Eq. (37) with Eq. (B) and using the residue theorem, we can deduce the coefficients
$C_{k}$ $=\underset{s=S_{k}}{\mathrm{Res}}\left(\dfrac{x(0)+b\Phi(s)}{s-a-be^{-s\tau}}\right)$
$=\dfrac{x(0)+b\Phi(S_{k})}{1+b\tau e^{-S_{k}\tau}}.$ (39)
Therefore, the exponential expansion of the solution can be written as
$x(t)=\sum_{k=-\infty}^{\infty}\dfrac{x(0)+b\Phi(S_{k})}{1+b\tau e^{-S_{k}\tau}}e^{S_{k}t}.$ (40)
Comparing Eqs. (40) and (36) with Eq. (32), we obtain the exponential expansion of the fundamental solution:
$x_{0}(t)=\sum_{k=-\infty}^{\infty}\dfrac{1}{1+b\tau e^{-S_{k}\tau}}e^{S_{k}t}.$ (41)